OpenStudy (anonymous):

Find the derivative of f(x) = 6x + 2 at x = 1.

OpenStudy (anonymous):

@ganeshie8 Do i just square the x as the others

OpenStudy (danjs):

the derivative of 6x + 1 is a constant,

OpenStudy (danjs):

\[\frac{ d }{ dx }(6x + 2) = 6\]

OpenStudy (anonymous):

Then i plug in the 1

OpenStudy (danjs):

for all values of x, y ' = 6 It is a horizontal line. No plug in needed.

OpenStudy (anonymous):

Oh okay so the answer is 6 then, just ignore the 1?

OpenStudy (danjs):

the graph of the derivative is a horizontal line at y = 6 so for any x value you look at, the y value will be 6

OpenStudy (danjs):

in other words y = 0(x) + 6

OpenStudy (danjs):

y(1) = 0(1) + 6 = 6

OpenStudy (anonymous):

Oh okay so it is 6 since x is 0

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

no the slope is zero, it is a horizontal line at y = 6 y = m*x+6 Where m is the slope = 0 so you see for any x value you choose, in this case 1, y = 0*(1) +6

OpenStudy (anonymous):

Im confusd on what the answer is

OpenStudy (danjs):

the answer is 6

OpenStudy (anonymous):

is it 1 or 6?

OpenStudy (danjs):

Do you see that if you have a horizontal line graphed at y = 6 no matter the x value, the y value is 6

OpenStudy (anonymous):

So any question in that form the answer is the y vallue?

OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -8 - 9t. Find the instantaneous velocity at t = 1 by finding the derivative. The answer would be -8 since it is thte y value?

OpenStudy (danjs):

you need to find the derivative first. That is the important part. In this case, the derivative is just a constant number 6.

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

\[\frac{ d }{ dt }[-8-9t]\]

OpenStudy (anonymous):

then do the y=mx+b

OpenStudy (danjs):

what is the derivative first?

OpenStudy (anonymous):

im honestly confused ,what is the equation

OpenStudy (danjs):

\[\frac{ d }{ dx }x ^{n}=n*x ^{n-1}\]

OpenStudy (danjs):

The power rule is what you would use

OpenStudy (danjs):

The derivative of [-8-9t] is -9 so for t=1, the instantaneous velocity is -9

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

thank you!!!!!!